OpenStudy (anonymous):

IF the length of one base of a trapezoid is 5, the length of the other base is 2x+7, and the length of the midsegment is 6x-24, what is the value of x? show all work that leads to your answer. Help Please:)

OpenStudy (anonymous):

@Mertsj @SolomonZelman @Loser66 Help Guys pleasee lol

OpenStudy (mertsj):

The midsegment is the average of the two bases.

OpenStudy (anonymous):

So I could devide by one of the bases to find the other?

OpenStudy (mertsj):

Do you know how to find an average?

OpenStudy (anonymous):

Not really

OpenStudy (mertsj):

What is the average of 6 and 4?

OpenStudy (anonymous):

5

OpenStudy (mertsj):

How did you get that?

OpenStudy (anonymous):

(6+4)/2

OpenStudy (mertsj):

Do that with the 2 bases of the trapezoid

OpenStudy (anonymous):

so (5+2x+7)/2?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

My calculator says 10.89

OpenStudy (mertsj):

So I guess you skipped school the day they had the lesson about "only like terms can be added"?

OpenStudy (anonymous):

Oh lol, 13+2x/2 then lol

OpenStudy (mertsj):

7+5=12

OpenStudy (anonymous):

Yes, yes it is lol

OpenStudy (mertsj):

\[\frac{2x+12}{2}\] Simplify that

OpenStudy (anonymous):

1x+6?

OpenStudy (mertsj):

Yes. That is the average of the two bases.

OpenStudy (anonymous):

Would that be my answer?

OpenStudy (mertsj):

The midsegment is the average of the two bases.

OpenStudy (mertsj):

What does the problem say the midsegment is?

OpenStudy (anonymous):

6x-24

OpenStudy (mertsj):

And what is the average of the two bases?

OpenStudy (anonymous):

1x+6

OpenStudy (mertsj):

The average of the two bases is equal to the midsegment. Write that equation.

OpenStudy (anonymous):

6x-24=1x+6

OpenStudy (mertsj):

Solve that equation

OpenStudy (anonymous):

6x=1x+30 5x=30 x=6

OpenStudy (mertsj):

Now answer the question the problem asks.

OpenStudy (anonymous):

Thank You Again @Mertsj

OpenStudy (mertsj):

yw